Harry Company sells 40,000 units at $33 per unit. Variable costs are $22.77 per unit, and fixed costs are $241,400.

Determine (a) the contribution margin ratio, (b) the unit contribution margin, and (c) income from operations.

a. Contribution margin ratio (Enter as a whole number.) %
b. Unit contribution margin (Round to the nearest cent.) $ per unit
c. Income from operations $
Answer #1

Contribution margin=Sales-Variable cost

=(33-22.77)=$10.23 per unit

Contribution margin ratio=Contribution margin/Sales

=10.23/33=31%

Total Contribution margin=(10.23*40,000)=$409200

Less:Fixed cost=(241,400)

Income from operations=$167800

Contribution margin ratio 31%
Unit Contribution margin $10.23
Income from operations $167800